# Influence of catastrophes and hidden dynamical symmetries on ultrafast backscattered photoelectrons

### v1

@@ V Conclusions
A key result of the present paper is that the maximal scattering angle is strongly influenced by the potential softening. While, for a hard-core Coulomb potential, the maximum rescattering angle is $\pi$ and the ridges close, for a soft-core potential the maximum rescattering angle is smaller and the ridges fold. This angle differs for specific pairs or orbits, and softening parameters, which suggests that the rescattering ridge is not purely kinematic. These findings are interpreted using catastrophe theory, which shows that the long and short backscattered orbits within a pair, as functions of the electron’s initial conditions, form a surface that folds where the two orbits coalesce, that is, at the ridge. Introducing a softening parameter creates two additional sets of solutions, which are qualitatively different. These solutions never touch the polarization axis, but have relatively low initial transverse momentum components. This implies that they will interact strongly with the core.
@@
Finally, our results show that, in the soft-core case, the surface defined by the minimal angle in the soft-core case acts as some kind of cutoff, although it depends on the interplay between the binding potential and the laser field. This indicates once more that, in a Coulomb-distorted scenario, it is hard to distinguish between dynamic and kinematic constraints. The present work revisits a well-known feature in laser-induced rescattering and brings a fresh perspective on the limitations of soft-core models. This is important due to the widespread use of soft-core potentials in the modeling of strong-field phenomena, and their resulting from laser-induced rescattering.
@@ Acknowledgements.
We thank A.S. Maxwell and B.B. Augstein for useful discussions. This work was funded by grants No. EP/J019143/1 and EP/T517793/1, from the UK Engineering and Physical Sciences Research Council (EPSRC).

### v2

@@ V Conclusions
A key result of the present paper is that the maximal scattering angle is strongly influenced by the potential softening. While, for a hard-core Coulomb potential, the maximum rescattering angle is $\pi$ and the ridges close, for a soft-core potential the maximum rescattering angle is smaller and the ridges fold. This angle differs for specific pairs or orbits, and softening parameters, which suggests that the rescattering ridge is not purely kinematic. These findings are interpreted using catastrophe theory, which shows that the long and short backscattered orbits within a pair, as functions of the electron’s initial conditions, form a surface that folds where the two orbits coalesce, that is, at the ridge. Introducing a softening parameter creates two additional sets of solutions, which are qualitatively different. These solutions never touch the polarization axis, but have relatively low initial transverse momentum components. This implies that they will interact strongly with the core. This effect is also present for ab-initio computations, but it manifests itself more subtly.
@@
For simplicity, our computations have been performed for linearly polarized monochromatic fields, but our findings hold for other driving-field scenarios, as long as recollision with the core is appreciable. Nonetheless, ultrashort pulses would add more complexity to the problem, as different half cycles would no longer be identical. This would influence the rescattering ridges and the PMDs’ centers [71, 72]. The unequal cycles would also lead to dominant events strongly dependent on the field parameters, and different field gradients for each half cycle [73, 74, 75]. Still, the same arguments could be applied in the few-cycle case, with the analysis performed for each ionization event separately. This can be understood by the fact that, upon rescattering, the field dresses the system via the vector potential, but the atomic potential dictates the dynamics. Furthermore, monochromatic waves are a good approximation for long enough pulses, and have successfully reproduced experimental findings in the context of the CQSFA. This holds even for very subtle features such as multi-path holographic interference [49, 41].
Moreover, we have verified that the folding due to the symmetry breaking reported in our paper for the soft-core potential also holds for longer wavelengths. This is important, as it is often assumed that the residual potentials are much less relevant in the longer wavelength regime. However, our studies show that the minimal scattering angle that occurs when softening is incorporated increases with the driving-field wavelength. Thus, even in the long-wavelength regime, the rescattering ridge is not a purely kinematic feature. This apparent counterintuitive result is caused by the the initial momentum maps being much more localized along the polarization axis. On the one hand, this localization means that the long-range potential tail loses importance, which is often used as a justification for employing Coulomb-free methods such as the Strong-Field Approximation. Nonetheless, this also means that, upon return, a backscattered electron will get much closer to the core than for near-IR driving fields. Consequently, the non-Coulomb behavior will be felt more strongly as the driving-field wavelength increases. An example of this momentum map is provided in the appendix.
Finally, our results show that, in the soft-core case, the surface defined by the minimal angle acts as some kind of cutoff, although it depends on the interplay between the binding potential and the laser field. This indicates once more that, in a Coulomb-distorted scenario, it is hard to distinguish between dynamic and kinematic constraints. The present work revisits a well-known feature in laser-induced rescattering and brings a fresh perspective on the limitations of soft-core models. Early studies for reduced-dimensionality models have meticulously investigated several properties of the soft-core potential, and concluded that for practical purposes it behaves like a Coulomb potential [76]. These studies encompass static properties such as scaling, energy levels, asymptotic behavior, threshold behavior, or the existence of a Rydberg-type series, and dynamic properties in the context of weak external fields. However, nowhere in this early paper it is mentioned how an electron scatters with this potential. Several decades later, our results show that there are key differences regarding the maximal scattering angles, which will reflect on the resulting photoelectron momentum distributions. This is important due to the widespread use of soft-core potentials in the modeling of strong-field phenomena, and their resulting from laser-induced rescattering.
@@ Acknowledgements.
We thank A.S. Maxwell and B.B. Augstein for useful discussions. This work was funded by grants No.EP/T019530/1 and EP/T517793/1, from the UK Engineering and Physical Sciences Research Council (EPSRC).
Appendix A Wavelength dependence
In this appendix, we illustrate how the folding effect discussed in the main body of this article persists and even becomes more pronounced for driving fields with longer wavelengths. These wavelengths have received a great deal of attention due to them unambiguously representing the quasi-static ionization regime [77]. Furthermore, the strong-field approximation is expected to work better in this regime.
Figure 12: Photoelectron momentum distributions, showing just the contribution of the earliest rescattered trajectories, for a range of softening parameter and wavelength with the remaining field parameters kept identical to those outlined in Fig. 2. The PMDs in panels (a) and (b) ((c) and (d)) were calculated for a 1300nm (2000nm) field with a softening parameter of $\alpha=1\times 10^{-6}$ and $\alpha=1\times 10^{-2}$ respectively. The white arrow on the PMDs indicates the maximum photoelectron energy, which occurs at an angle $\theta$ to the $p_{f,\parallel}$ axis.
Figure 13: Momentum mapping of the earliest rescattering event for the 2000nm field with softening parameter $1\times 10^{-2}$ and all other parameters as described in Fig. 2. The dashed coloured lines and other aspects of the figure have meanings analogous to those in Fig. 7. Panels (a) and (b) give the final and initial momenta, respectively.
Fig. 12 displays the PMDs obtained for the hard-core [Figs. 12(a) and (c)] and soft-core [Figs. 12(b) and (d)] potential, calculated with only the shortest pair of backscattered trajectories using the wavelengths of $\lambda=1300$ nm and $\lambda=2000$ nm (first and last two rows from the top, respectively). The angles found in the figure correspond to the scatter points in Fig. 11, in the main body of the paper.
The figure shows that the ridges do not close for the soft-core case, and that the minimal rescattering angle increases with $\lambda$. This is a counterintuitive result, which illustrates that care must be taken when using softening in the long wavelength regime. The same trend is observed in the approximate scan over the wavelengths shown in Fig. 11, computed with our analytical model.
The initial to final momentum mapping for the 2000 nm field and a softening parameter of $\alpha=1\times 10^{-2}$ sheds light on this behavior. This mapping is presented in Fig. 13 which corresponds to the PMD shown in Fig. 12(d). The four sheets encountered for the $\lambda=800$ nm field are once more present, with the final momenta being restricted to specific angular regions [see Figs. 13(a)]. However, the initial momentum mapping, displayed in Fig. 13(b), is much more localized along the polarization axis than in the $\lambda=800$ nm case. This is expected, as the Coulomb tail will lose relevance for $\lambda=800$ nm. On the other hand, a momentum distribution located around vanishingly small perpendicular momenta means that the interaction with the core will be stronger and that the electron will assess a region in which the scattering properties are markedly different for the hard- and soft-core potentials. This implies that the non-Coulomb behavior arising from the soft-core potential will play a more important role as the wavelength increases.